OpenStudy (anonymous):

what is the domain of this ? x^2 over sqrt{6-x}

OpenStudy (anonymous):

\[f(x)=x^2/\sqrt{6-x}\]\[D_f=\left\{ xεR/6-x >0 \right\}\]\[D_f=\left\{ xεR/x <6 \right\}\]\[D_f=(-\infty,6)\]